OpenStudy (anonymous):

Find an equation of the line that passes through the given points. (Let x be the independent variable and y be the dependent variable.) (2, 8) and (3, 15)

OpenStudy (anonymous):

Okay first find the gradient/slope: (15-8)/(3-2) = 7/1= 7 Eqn of line y= 7x +c where c is a constant, now we got to find c Substitute any point, (2,8) into the eqn. 8 = 7*2 +c 8= 14 +c c = -6 Thus the eqn is y = 7x -6

OpenStudy (anonymous):

wow this was really simple i made it so difficult. thank you!